chocolateluva77:

can smo reply asap? dis is rationalizing denominators: index 3 or higher how do u simplify dis? pls explain how u got the steps to the solution ty

chocolateluva77:

Tranquility:

When you rationalize the denominator, you are trying to get rid of the cube root of the denominator. You need to multiply the numerator and denominator by \(\sqrt[3]{225}\) twice so that way the cube root cancels out in the denominator Essentially, \(\dfrac{2}{\sqrt[3]{225}} \times \dfrac{\sqrt[3]{225}}{\sqrt[3]{225}} \times \dfrac{\sqrt[3]{225}}{\sqrt[3]{225}} = ?\)

chocolateluva77:

so what do u do after?

Tranquility:

Simplify it

chocolateluva77:

how do u rationalize the denominator index 3√225

Tranquility:

When you multiply the denominator across, you get \( \left(\sqrt[3]{225} \right)^3\) What can you simplify that to?

chocolateluva77:

1 bruh idk

Tranquility:

No... how?

Tranquility:

cube root is the same thing as to the power of 1/3 \(\sqrt[3]{x} = x^{\frac{1}{3}}\) Another way to write \((\sqrt[3]{225})^3\) would be as \( (225^{\frac{1}{3}})^3\) Can you simplify that now? All you have to do is multiply the exponents.

chocolateluva77:

is it (676/3) exponent 3?

Tranquility:

no

Tranquility:

\((\sqrt[3]{225})^3\) = \( (225^{\frac{1}{3}})^3 = 225 ^{(\frac{1}{3} \times 3)} = ??\)

chocolateluva77:

225

Tranquility:

Exactly

Tranquility:

\(\dfrac{2}{\sqrt[3]{225}} \times \dfrac{\sqrt[3]{225}}{\sqrt[3]{225}} \times \dfrac{\sqrt[3]{225}}{\sqrt[3]{225}} = ?\) \( = \dfrac{2 \times \sqrt[3]{225} \times \sqrt[3]{225}}{225}\) Can you simplify the numerator? I'll help start you off \( = \dfrac{2 \times \sqrt[3]{225 \times 225}}{225}\) Remember that 225 is equal to 15 * 15 So we have 15 * 15 * 15 * 15 under the cube root. Can you factor out 15 from it?

Tranquility:

Basically, you need to use this property: \(\large\sqrt[3]{x^3} = x\)

chocolateluva77:

i don't get it 😭 how do i factor out 15?

Tranquility:

We're only looking at the cube root right now \( \sqrt[3]{225\times 225} = \sqrt[3]{15^3 \times 15} = \sqrt[3]{15^3} \times \sqrt[3]{15}\) Do you see it now?

chocolateluva77:

why did u do 15 exponent 3 × 15?

chocolateluva77:

15 exponent 3 is not 225

Tranquility:

225 is 15 * 15 225 * 225 is going to be 15*15*15*15 which is 15^3 * 15

chocolateluva77:

so the answer is 2 index 3 √15/15

chocolateluva77:

dis is sm steps man is there a easier method?

chocolateluva77:

wya?

Tranquility:

@chocolateluva77 wrote:
so the answer is 2 index 3 √15/15
Yes! \(\dfrac{2\sqrt[3]{15}}{15}\)